OpenStudy (anonymous):

whats the correct simplification of (4x3y4z5)(5x4y5z3).

OpenStudy (anonymous):

\[\huge (4x^3y^4z^5)(5x^4y^5z^3)\] is this what you meant?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\LARGE x^a\cdot x^b=x^{a+b}\] you're supposed to know this rule... do you know it ?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

well, you have to use it, solving this ! example: \[\LARGE 3^4\cdot 3^7=3^{4+7}\] the same with: \[\LARGE z^2\cdot z^9=z^{2+9}\] now you just take off bracket.. you don't need them ! and multiply by following these examples and you should end up with a solution which is... \[\LARGE 20x^7z^8y^9\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Glad to help :)